OpenStudy (anonymous):

Solve equation: xy + x+y = 71 x^2*y + x*y^2=880 Found x and y and then calculate for x^2 + y^2

OpenStudy (aaronandyson):

xy + x + y = 71 x^2y + xy^2 = 880 using 2nd eq then xy(x+y)=880 then using 1st eq xy(71-xy)=880 let w=xy then -w^2+71w-880=0 xy=16 or xy=55 and x+y=55 or x+y=16 case 1)xy=16 and x+y=55 (is not possible) because x and y positive integers case 2)xy=55 and x+y=16 (x=11,y=5 or x=5,y=11) x^2+y^2= (x+y)^2-2xy = 16^2-2(55)=146 Answer: x^2+y^2=146

OpenStudy (anonymous):

Thanks!

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :3